# A Risk Taxonomy for Evaluating AI-Powered Psychotherapy Agents

### v1

@@ Abstract.
The proliferation of Large Language Models (LLMs) and Intelligent Virtual Agents acting as psychotherapists presents significant opportunities for expanding mental healthcare access. However, their deployment has also been linked to serious adverse outcomes, including user harm and suicide, facilitated by a lack of standardized evaluation methodologies capable of capturing the nuanced risks of therapeutic interaction. Current evaluation techniques lack the sensitivity to detect subtle changes in patient cognition and behavior during therapy sessions that may lead to subsequent decompensation. We introduce a novel risk taxonomy specifically designed for the systematic evaluation of conversational AI psychotherapists. Developed through an iterative process including review of the psychotherapy risk literature, qualitative interviews with clinical and legal experts, and alignment with established clinical criteria (e.g., DSM-5) and existing assessment tools (e.g., NEQ, UE-ATR), the taxonomy aims to provide a structured approach to identifying and assessing user/patient harms. We provide a high-level overview of this taxonomy, detailing its grounding, and discuss potential use cases. We discuss two use cases in detail: monitoring cognitive model-based risk factors during a counseling conversation to detect unsafe deviations, in both human-AI counseling sessions and in automated benchmarking of AI psychotherapists with simulated patients. The proposed taxonomy offers a foundational step towards establishing safer and more responsible innovation in the domain of AI-driven mental health support.
AI Psychotherapy, Risk Assessment, Intelligent Virtual Agents, LLM Safety, Mental Health Technology
journalyear: 2025conference: ; May 2025; booktitle: Preprintccs: Applied computing Consumer healthccs: Information systems Language modelsccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Artificial Intelligence (AI) systems, particularly those leveraging Large Language Models (LLMs), are increasingly being developed and deployed as tools for mental health support, including entertainment chatbots ”masquerading” as therapists (Jurblum and Selzer, 2025; Barry, 2025; Jin et al., 2025; Na et al., 2025). The American Psychological Association has recently warned federal regulators about AI chatbots programmed to reinforce, rather than challenge, a user’s thinking, potentially driving vulnerable users to harm themselves or others (Lawrence et al., 2024; Barry, 2025). Reports have emerged, including instances where interactions with AI chatbot ”therapists” have been linked to severe negative outcomes. Notably, one case involved a 14-year-old boy’s suicide after months of interaction with chatbots on Character.AI, including a chatbot that falsely claimed to have several years’ experience as a licensed therapist (Roose, 2024; Barry, 2025). In this instance, the chabot reportedly reinforced the user’s harmful ideations rather than challenging them (Barry, 2025). Such events underscore the profound ethical responsibilities and potential harms associated with deploying these technologies to vulnerable populations, often with concerns raised about them being ”dangerous and untested” or lacking rigorous, context-appropriate safety evaluations (Roose, 2024; Barry, 2025; Stade et al., 2024b).
A critical issue is the absence of standardized methods to evaluate these AI systems for the unique risks inherent in psychotherapeutic interactions. Existing LLM benchmarks often focus on general capabilities such as toxicity detection in isolated texts, failing to capture harms that may emerge from sustained, dynamic conversations (Chang et al., 2023; Zhang et al., 2025). Traditional clinical trials, while vital for efficacy, can be slow and ethically challenging for initial safety validation of unproven autonomous systems (Stade et al., 2024b; Heinz et al., 2025). This gap leaves users susceptible to unexpected adverse outcomes.
To address this, we propose a risk taxonomy tailored for evaluating AI psychotherapists. We present a high-level overview of this taxonomy, derived from an iterative process including a review of the literature on psychotherapy risks, qualitative interviews with clinical and legal experts, and alignment with established clinical criteria and existing assessment tools. We hypothesize that this taxonomy can provide a structured and systematic approach to identifying, categorizing, and ultimately mitigating risks in AI-driven psychotherapy applications. Our primary aim is to provide a framework that can be used to evaluate automated psychotherapy systems for potential negative impacts before and during their deployment, fostering safer innovation in digital mental health.
@@ 2.1. Risks in Psychotherapy
Several frameworks attempt to categorize negative experiences in psychotherapy. Linden and Schermuly-Haupt distinguish Unwanted Events (UEs) from Adverse Treatment Reactions (ATRs) caused by correctly applied treatment, and malpractice reactions from incorrectly applied treatment (Linden, 2013; Linden and Schermuly-Haupt, 2014). Mejía-Castrejón et al. (Mejía-Castrejón et al., 2024) adapted a medical model, classifying AEs by severity, relatedness, seriousness, and expectedness. Common AEs identified across studies and captured in assessment tools like the Negative Effects Questionnaire (NEQ) (Rozental et al., 2019) or the Unwanted Event to Adverse Treatment Reaction (UE-ATR) checklist (Linden, 2013) include negative emotions (e.g., anxiety, sadness), symptom worsening, the emergence of unpleasant memories, relationship strains, dependency on therapy, or reduced self-efficacy (Linden and Schermuly-Haupt, 2014; Yao et al., 2020). Factors influencing these include patient and therapist characteristics, and specific techniques (Linden and Schermuly-Haupt, 2014).
Assessing these effects is difficult; therapists may struggle to recognize negative outcomes, and establishing causality is complex (Linden, 2013; Linden and Schermuly-Haupt, 2014). While tools like the UE-ATR, NEQ, and the Inventory for Negative Effects of Psychotherapy (INEP) (Ladwig et al., 2014) exist, consensus on definitions and instruments remains elusive (Mejía-Castrejón et al., 2024). Reviews show variability and inadequate reporting of negative effects in randomized clinical trials (Jonsson et al., 2014; Klatte et al., 2023), underscoring the need for standardized assessment. Our taxonomy development aims to build upon these existing efforts by providing a structured framework specially adapted for evaluating AI-powered psychotherapy agents.
@@ 2.2. Safety Evaluation of LLMs in Healthcare
The application of LLMs in high-stakes domains like healthcare, including psychotherapy, demands rigorous safety assessment due to the potential for significant harm (Stade et al., 2024b; Lawrence et al., 2024; Steenstra et al., 2024). General LLM safety evaluations often focus on risks including bias, toxicity, or misinformation (Weidinger et al., 2021; Ji et al., 2023), using benchmarks such as ALERT (Tedeschi et al., 2024) or SafetyBench (Zhang et al., 2023). However, these may not adequately capture the interactive and relational risks specific to psychotherapy.
@@
Safety frameworks are emerging for AI in mental health, including the READI framework (Stade et al., 2024a) and the BPES checklist, emphasizing ”Do No Harm” (Khan and Seto, 2023). Dialogue safety in mental health support is also a focus (Qiu et al., 2023). For instance, $\Psi$-ARENA offers an interactive platform for evaluating LLM-based psychological counselors (Zhu et al., 2025). Despite these advances, a standardized taxonomy to systematically identify and categorize the spectrum of potential risks specific to AI psychotherapeutic interactions, grounded in clinical realities and existing harm assessment principles, is still lacking. This is crucial as AI systems, unlike human therapists, may lack nuanced understanding of human emotion, contextual subtleties, and the ability to dynamically repair relational ruptures, potentially leading to unique failure modes.
3. Taxonomy Development
@@
The development of our risk taxonomy followed an iterative design process. We first formulated a preliminary taxonomy based on an in-depth literature review of psychotherapy risk (e.g., (Linden, 2013; Rozental et al., 2018; Klatte et al., 2023)). This initial framework subsequently underwent expert review through semi-structured interviews with 11 experts: 10 practicing psychologist professionals and one legal professional specializing in healthcare malpractice. These interviews were designed to explore perceptions of risk in human therapy, identify potential failure modes for AI therapists, and gather direct feedback on the preliminary taxonomy. The interview study was approved by Northeastern University’s IRB.
The interviews were analyzed using thematic analysis (Clarke and Braun, 2017) to reveal six key themes.
3.1. Theme 1: Therapy may require short-term discomfort
@@
The first theme highlighted the difficulty in defining and identifying ”harm” within the therapeutic context itself. Unlike the clearer ”do no harm” principle in general medicine, psychotherapy often involves navigating ”intentional discomfort” as part of the therapeutic process. Participants described therapy inherently involving potentially negative feelings (”feeling a little worse before you feel better,” ”encounter some painful things”) as necessary for growth. Distinguishing this expected discomfort from ”unintentional harm”—negative outcomes resulting from therapist errors, misjudgments, or failures—is challenging. Harm exists on a ”spectrum,” ranging from temporary setbacks (”bruises”) like feeling ”misunderstood” or ”judged,” to more significant consequences (”cuts”) such as a ”rupture” in the ”therapeutic alliance,” premature termination (”stop coming to therapy,” ”drop out early”), worsening symptoms (”exacerbate symptoms”), loss of ”trust in self,” developing a ”bad taste… about therapy,” or, in the most severe cases, increased ”suicide” risk. Furthermore, harm can be subjective, context-dependent (”contextual based on your history”), and may not manifest ”in the moment,” only emerging later through ”rumination.”
3.2. Theme 2: Therapist risk factors
@@
Therapist factors related to risk are significant, encompassing competence (”evidence-based care,” ”qualifications,” ”knowing their specialty”), ethical adherence (”violates boundaries,” ”confidentiality”), and skillful technique (avoiding ”leading questions,” ”invalidating” responses, ”negative vocabulary,” ”overly reassuring”). Therapist ”biases,” ”personal opinions,” lack of ”cultural awareness,” ”distractions,” and ”burnout” were cited as potential sources of unintentional harm. Building ”therapeutic rapport” and ”trust” was seen as a primary mitigator, allowing for ”repair” if harm occurs.
3.3. Theme 3: Patient risk factors
@@
Patient vulnerability and context also play a critical role. ”Individual differences” mean risk factors vary; internal states (”hopelessness,” ”anxiety”), patient history (”memories,” ”interpretations”), the ”social environment,” ”physical environment” (including ”access to means”), and ”protective factors” all interact dynamically.
3.4. Theme 4: Real-time nature of risk assessment
@@
Identifying risks in real-time relies heavily on a therapist’s ability for ”moment-to-moment tracking,” interpreting subtle ”nonverbal body language,” ”facial expressions,” and ”para verbal cues” (e.g., ”intonation”), alongside verbal content and formal ”assessment tools” (e.g., ”Columbia Scale”(Posner et al., 2008), ”PHQ-9”(Kroenke et al., 2001)). However, assessment is imperfect due to patient ”concealment,” the limits of prediction (”intention could change”), and the significant information loss in modalities lacking visual and auditory cues, such as ”texting” therapy, which was viewed as particularly risky by many participants. Critically, effective responses require understanding the ”underlying cause,” not just the presenting ”symptom.”
3.5. Theme 5: Perceived shortcomings of AI therapists
@@
When considering AI therapists, participants largely believed the fundamental risks in human therapy remain the same but are potentially ”higher” or exacerbated. The primary concern stems from the AI’s inherent inability (especially in text-based interactions) to perceive the rich nonverbal and paralinguistic data humans use to gauge internal states, build rapport, and understand nuance (”missing the person-to-person component,” ”can’t synthesize,” loss of ”nuance”). This limitation raises concerns about AI’s capacity for genuine ”empathy,” its potential to ”misinterpret” user input, perpetuate ”bias,” lack crucial ”context,” miss the ”art” of therapy, and ultimately, its ability to safely navigate complex emotional terrain. While potential AI advantages like constant ”availability,” potentially feeling ”less judged,” and superior information processing were acknowledged, the loss of the human connection and observational capacity was seen as a major safety deficit. Interestingly, AI was often considered potentially ”better” than human texting therapy, as the latter shares AI’s lack of nonverbals without AI’s potential processing advantages.
3.6. Theme 6: Refinement of risk taxonomy
@@
The feedback on the proposed risk taxonomy indicated it was a ”good start” covering many ”big risks,” but highlighted needs for refinement. Key omissions noted were ”trauma” and ”abuse.” Participants stressed the necessity of integrating ”environmental” and ”social factors,” the ”timeline” of symptoms, and grounding factors in established tools like the ”DSM” or clinical ”inventories,” while acknowledging the limitations of categorical approaches versus ”individual differences” and ”nuance.” The taxonomy’s current structure was seen as potentially useful but insufficient alone for risk assessment without incorporating deeper context and validated measures.
These insights drove significant revisions: patient factors within the risk taxonomy were linked to the DSM-5, factor values are now compared to baselines to accommodate individual differences, and risk categorization was grounded in approaches similar to the NEQ and UE-ATR.
4. Overview of the Risk Taxonomy

### v2

# A Risk Ontology for Evaluating AI-Powered Psychotherapy Virtual Agents

@@ Abstract.
The proliferation of Large Language Models (LLMs) and Intelligent Virtual Agents acting as psychotherapists presents significant opportunities for expanding mental healthcare access. However, their deployment has also been linked to serious adverse outcomes, including user harm and suicide, facilitated by a lack of standardized evaluation methodologies capable of capturing the nuanced risks of therapeutic interaction. Current evaluation techniques lack the sensitivity to detect subtle changes in patient cognition and behavior during therapy sessions that may lead to subsequent decompensation. We introduce a novel risk ontology specifically designed for the systematic evaluation of conversational AI psychotherapists. Developed through an iterative process including review of the psychotherapy risk literature, qualitative interviews with clinical and legal experts, and alignment with established clinical criteria (e.g., DSM-5) and existing assessment tools (e.g., NEQ, UE-ATR), the ontology aims to provide a structured approach to identifying and assessing user/patient harms. We provide a high-level overview of this ontology, detailing its grounding, and discuss potential use cases. We discuss four use cases in detail: monitoring real user interactions, evaluation with simulated patients, benchmarking and comparative analysis, and identifying unexpected outcomes. The proposed ontology offers a foundational step towards establishing safer and more responsible innovation in the domain of AI-driven mental health support.
AI Psychotherapy, Risk Assessment, Intelligent Virtual Agents, LLM Safety, AI Alignment, Mental Health Technology
journalyear: 2025conference: ; September 2025; booktitle: Preprintdoi: 10.1145/3717511.3749286isbn: 979-8-4007-1508-2/2025/09ccs: Applied computing Consumer healthccs: Information systems Language modelsccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Intelligent Virtual Agents (IVAs), particularly those powered by Large Language Models (LLMs), are increasingly deployed as AI-powered psychotherapy agents, sometimes by entertainment chatbots ”masquerading” as qualified professionals (Jurblum and Selzer, 2025; Barry, 2025; Jin et al., 2025; Na et al., 2025). This trend raises serious safety concerns, highlighted by the American Psychological Association’s warning about AI reinforcing harmful thinking (Barry, 2025). Alarming incidents, such as a user’s suicide linked to interactions with an AI chatbot ”therapist” that reportedly amplified harmful ideations (Roose, 2024; Barry, 2025), underscore the profound risks. These events reveal the dangers of deploying such AI-powered IVAs, often ”dangerous and untested,” to vulnerable individuals without rigorous, context-appropriate safety evaluations (Roose, 2024; Barry, 2025; Stade et al., 2024).
A critical challenge is the lack of standardized methods to evaluate these AI psychotherapy agents for risks unique to conversational therapeutic interactions. Current LLM benchmarks assess general capabilities like toxicity, but are ill-suited for the sustained, dynamic conversations inherent when AI-powered IVAs conduct therapy (Chang et al., 2024; Zhang et al., 2025). Traditional clinical trials, vital for efficacy, are also slow and ethically challenging for initial safety validation of autonomous AI psychotherapists (Stade et al., 2024; Heinz et al., 2025). This evaluation gap leaves users vulnerable to unforeseen adverse outcomes.
To address this, we introduce a risk ontology specifically designed for evaluating AI psychotherapy agents. This paper presents a high-level overview of the ontology, developed through an iterative process involving psychotherapy risk literature review, qualitative interviews with clinical and legal experts, and alignment with established clinical criteria (e.g., DSM-5) and assessment tools (e.g., NEQ, UE-ATR). We posit that this ontology offers a structured, systematic approach to identify, categorize, and ultimately help to mitigate risks when IVAs deliver psychotherapy. Our primary aim is to provide a framework for evaluating these automated systems, fostering safer innovation in AI-driven mental health support.
@@ 2.1. Risks in Psychotherapy
Several frameworks attempt to categorize negative experiences in psychotherapy. Linden and Schermuly-Haupt distinguish Unwanted Events (UEs) from Adverse Treatment Reactions (ATRs) caused by correctly applied treatment, and malpractice reactions from incorrectly applied treatment (Linden, 2013; Linden and Schermuly-Haupt, 2014). Mejía-Castrejón et al. (Mejía-Castrejón et al., 2024) adapted a medical model, classifying AEs by severity, relatedness, seriousness, and expectedness. Common AEs identified across studies and captured in assessment tools like the Negative Effects Questionnaire (NEQ) (Rozental et al., 2016) or the Unwanted Event to Adverse Treatment Reaction (UE-ATR) checklist (Linden, 2013) include negative emotions (e.g., anxiety, sadness), symptom worsening, the emergence of unpleasant memories, relationship strains, dependency on therapy, or reduced self-efficacy (Linden and Schermuly-Haupt, 2014; Yao et al., 2020). Factors influencing these include patient and therapist characteristics, and specific techniques (Linden and Schermuly-Haupt, 2014).
Assessing these effects is difficult; therapists may struggle to recognize negative outcomes, and establishing causality is complex (Linden, 2013; Linden and Schermuly-Haupt, 2014). While tools like the UE-ATR, NEQ, and the Inventory for Negative Effects of Psychotherapy (INEP) (Ladwig et al., 2014) exist, consensus on definitions and instruments remains elusive (Mejía-Castrejón et al., 2024). Reviews show variability and inadequate reporting of negative effects in randomized clinical trials (Jonsson et al., 2014; Klatte et al., 2023), underscoring the need for standardized assessment. Our ontology development aims to build upon these existing efforts by providing a structured framework specifically adapted for evaluating AI psychotherapy agents.
@@ 2.2. Safety Evaluation of LLMs in Healthcare
The application of LLMs in high-stakes domains like healthcare, including psychotherapy, demands rigorous safety assessment due to the potential for significant harm (Stade et al., 2024; Lawrence et al., 2024; Steenstra et al., 2024). General LLM safety evaluations often focus on risks including bias, toxicity, or misinformation (Weidinger et al., 2022; Ji et al., 2025), using benchmarks such as ALERT (Tedeschi et al., 2024) or SafetyBench (Zhang et al., 2024). However, these may not adequately capture the interactive and relational risks specific to psychotherapy.
@@
Safety frameworks are emerging for AI in mental health, including the READI framework (Stade et al., 2025) and the BPES checklist, emphasizing ”Do No Harm” (Khan and Seto, 2023). Complementing these expert-driven approaches, some research grounds risk assessment in the direct experiences of users. For instance, Chandra et al. (Chandra et al., 2025) developed a psychological risk taxonomy by surveying 283 individuals with lived mental health experience about their negative interactions with AI conversational agents, categorizing risks into specific AI behaviors, negative psychological impacts, and the user contexts that mediate them. Dialogue safety in mental health support is also a focus (Qiu et al., 2023). For instance, $\Psi$-ARENA offers an interactive platform for evaluating LLM-based psychological counselors (Zhu et al., 2025). Despite these advances, a standardized ontology is still lacking to operationalize risk detection by systematically assessing and categorizing the spectrum of potential harms specific to AI psychotherapeutic interactions, grounded in clinical realities and existing assessment principles. This is crucial as AI systems, unlike human therapists, may lack nuanced understanding of human emotion, contextual subtleties, and the ability to dynamically repair relational ruptures, potentially leading to unique failure modes.
3. Ontology Development
@@
The development of our risk ontology followed an iterative design process. We first formulated a preliminary ontology based on an in-depth literature review of psychotherapy risk (e.g., (Linden, 2013; Rozental et al., 2018, 2016; Klatte et al., 2023)). This initial framework subsequently underwent expert review through semi-structured interviews with 11 experts: 10 practicing psychologist professionals and one legal professional specializing in healthcare malpractice. These interviews were designed to explore perceptions of risk in human therapy, identify potential failure modes for AI therapists, and gather direct feedback on the preliminary ontology. The interview study was approved by Northeastern University’s IRB. The interviews were analyzed using thematic analysis (Clarke and Braun, 2017) to reveal six key themes that directly informed the final structure and content of the ontology. The interview study was approved by Northeastern University’s IRB.
@@
Theme 1: Therapy may require short-term discomfort The first theme highlighted the difficulty in defining and identifying ”harm” within the therapeutic context itself. Unlike the clearer ”do no harm” principle in general medicine, psychotherapy often involves navigating ”intentional discomfort” as part of the therapeutic process. Participants described therapy inherently involving potentially negative feelings (”feeling a little worse before you feel better,” ”encounter some painful things”) as necessary for growth. Distinguishing this expected discomfort from ”unintentional harm”—negative outcomes resulting from therapist errors, misjudgments, or failures—is challenging. Harm exists on a ”spectrum,” ranging from temporary setbacks (”bruises”) like feeling ”misunderstood” or ”judged,” to more significant consequences (”cuts”) such as a ”rupture” in the ”therapeutic alliance,” premature termination (”stop coming to therapy,” ”drop out early”), worsening symptoms (”exacerbate symptoms”), loss of ”trust in self,” developing a ”bad taste… about therapy,” or, in the most severe cases, increased ”suicide” risk. Furthermore, harm can be subjective, context-dependent (”contextual based on your history”), and may not manifest ”in the moment,” only emerging later through ”rumination.”
@@
Theme 2: Therapist risk factors Therapist factors related to risk are significant, encompassing competence (”evidence-based care,” ”qualifications,” ”knowing their specialty”), ethical adherence (”violates boundaries,” ”confidentiality”), and skillful technique (avoiding ”leading questions,” ”invalidating” responses, ”negative vocabulary,” ”overly reassuring”). Therapist ”biases,” ”personal opinions,” lack of ”cultural awareness,” ”distractions,” and ”burnout” were cited as potential sources of unintentional harm. Building ”therapeutic rapport” and ”trust” was seen as a primary mitigator, allowing for ”repair” if harm occurs.
@@
Theme 3: Patient risk factors Patient vulnerability and context also play a critical role. ”Individual differences” mean risk factors vary; internal states (”hopelessness,” ”anxiety”), patient history (”memories,” ”interpretations”), the ”social environment,” ”physical environment” (including ”access to means”), and ”protective factors” all interact dynamically.
@@
Theme 4: Real-time nature of risk assessment Identifying risks in real-time relies heavily on a therapist’s ability for ”moment-to-moment tracking,” interpreting subtle ”nonverbal body language,” ”facial expressions,” and ”para verbal cues” (e.g., ”intonation”), alongside verbal content and formal ”assessment tools” (e.g., ”Columbia Scale”(Posner et al., 2008), ”PHQ-9”(Kroenke et al., 2001)). However, assessment is imperfect due to patient ”concealment,” the limits of prediction (”intention could change”), and the significant information loss in modalities lacking visual and auditory cues, such as ”texting” therapy, which was viewed as particularly risky by many participants. Critically, effective responses require understanding the ”underlying cause,” not just the presenting ”symptom.”
@@
Theme 5: Perceived shortcomings of AI Psychotherapists When considering AI therapists, participants largely believed the fundamental risks in human therapy remain the same but are potentially ”higher” or exacerbated. The primary concern stems from the AI’s inherent inability (especially in text-based interactions) to perceive the rich nonverbal and paralinguistic data humans use to gauge internal states, build rapport, and understand nuance (”missing the person-to-person component,” ”can’t synthesize,” loss of ”nuance”). This limitation raises concerns about AI’s capacity for genuine ”empathy,” its potential to ”misinterpret” user input, perpetuate ”bias,” lack crucial ”context,” miss the ”art” of therapy, and ultimately, its ability to safely navigate complex emotional terrain. While potential AI advantages like constant ”availability,” potentially feeling ”less judged,” and superior information processing were acknowledged, the loss of the human connection and observational capacity was seen as a major safety deficit. Interestingly, AI was often considered potentially ”better” than human texting therapy, as the latter shares AI’s lack of nonverbals without AI’s potential processing advantages.
@@
Theme 6: Refinement of risk ontology The feedback on the proposed risk ontology indicated it was a ”good start” covering many ”big risks,” but highlighted needs for refinement. Key omissions noted were ”trauma” and ”abuse.” Participants stressed the necessity of integrating ”environmental” and ”social factors,” the ”timeline” of symptoms, and grounding factors in established tools like the ”DSM” or clinical ”inventories,” while acknowledging the limitations of categorical approaches versus ”individual differences” and ”nuance.” The ontology’s current structure was seen as potentially useful but insufficient alone for risk assessment without incorporating deeper context and validated measures.
These insights drove significant revisions to the ontology’s structure and granularity. Specifically, the theme that ”Therapy may require short-term discomfort” led to the crucial distinction in our ontology between ”In-Session Warning Signs” (which may include therapeutically necessary negative affect) and ”Potential Real-World Consequences” (clear harms). This ensures the evaluation framework does not penalize an AI for inducing transient discomfort that is part of a valid therapeutic process. Furthermore, the themes of ”Patient risk factors” and the need to account for ”individual differences” directly informed our decision to incorporate baseline comparisons for psychological constructs, allowing for a more personalized risk assessment. Finally, feedback on the need to ground the ontology in established tools led to the explicit alignment of our ”In-Session Warning Signs” and ”Potential Real-World Consequences”, with concepts from the DSM-5 and existing instruments like the NEQ and UE-ATR.
4. Overview of the Risk Ontology